# PQMass: Probabilistic Assessment of the Quality of Generative Models using Probability Mass Estimation

### v2

@@ 2.1 Testing equality of quantized distributions using samples
With this definition, (1) can be restated: $p=q$ if and only if $\pi_{*}p=\pi_{*}q$ for all partitions $\mathbb{R}^{d}=R_{1}\sqcup R_{2}$, or, more generally, $\mathbb{R}^{d}=R_{1}\sqcup\dots\sqcup R_{n_{R}}$ (note that $\pi$ implicitly depends on the partition). As we will return to in § 2.2, it is sometimes sufficient to consider restricted classes of partitions.
The main idea behind PQMass is that the probability that $p=q$ is upper-bounded by the probability that $\pi_{*}p=\pi_{*}q$ for a given partition. Next, we will consider how to test the hypothesis that $\pi_{*}p=\pi_{*}q$ using samples from $p$ and $q$ given a fixed partition. In § 2.2 we discuss the choice of partition and propose an effective method.
@@ 2.2 Choosing an effective partition
The sets $R_{1},\dots,R_{n_{R}}$ then partition $\mathbb{R}^{d}$ into nonoverlapping regions, and the function $\pi$ that maps points to the index of the containing region is defined by $\pi(x)=\operatorname*{arg\,min}_{j}\|x-z_{j}\|$, breaking ties in favour of the lower index (see Fig. 1, right).
While the definition (6) uses the $L^{2}$ distance metric, one can define Voronoi cells using any metric $D:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$. In Table 3 we compare this choice to that of using the $L^{1}$ metric, as well as with the choice of using a metric defined in a feature space.
@@
For the validity of the test in § 2.1, the choice of the regions should be independent of the points used to perform the test. Thus one can use a small subset of the available sample points as reference points to construct a tessellation, then evaluate the $\chi^{2}_{\rm PQM}$ statistic using the remaining points.
@@ 2.3 Consistency guarantees: What information is lost by PQMass?
Since $\pi$ depends on $r$, the probability referred to in Prop. 2.1 is over the choice of $n_{R}$ samples from $r$. Of course, for "bad" choices of reference points, we may be unlucky and not distinguish $p$ and $q$ (if the two resulting cells have the same mass under $p$ and $q$). The proposition states that with positive probability, this does not happen. Similarly, as the number of references points $n_{R}$ grows to infinity, the probability that the test distinguishes two distributions approaches one. This can be understood as a consequence of the fact that the Voronoi cells ‘tessellate’ the space more finely as $n_{R}$ grows.
@@
Note that the $n_{R}\to\infty$ limit in Prop. 2.2 does not refer to performing the test with "$n_{R}=\infty$" to compare two $\chi^{2}$ distributions with infinite degrees of freedom. The proposition states that the probability that the tessellation defined using $n_{R}$ random reference points sampled from $r$ separates $p$ and $q$ – which is a well-defined quantity for any $n_{R}$ – approaches 1 as $n_{R}\to\infty$. These results show that, at least for ‘well-behaved’ densities, no information is lost by the PQMass statistic. By this, we mean that if two distributions are distinct, then the test has a positive probability of distinguishing them (and the probability approaches 1 in the asymptotic limit). This contrasts with tests that approximate the distributions by a certain fixed class (e.g., tests considering only summary statistics such as the second-moment statistics used by FID). These results can likely be generalized further, which we leave for future work.
Since Prop. 2.2 is about the sensitivity of the test in the asymptotic limit, the proposition simply tells us that the sensitivity of the test approaches 1 as the number of samples grows, but does not tell us about the rate of convergence. The usefulness of the test itself is demonstrated by the empirical evidence in the experiments presented in the next section.
@@ 3.1 Null test
We start by validating PQMass on a null test, by comparing two sets of samples that are known to be equivalent. As shown in Fig. 2, we use a Gaussian mixture model in $100$ dimensions with $20$ components as our generative model. We then repeat the following process $2^{14}$ times: We generate a number of samples (see legend) from the Gaussian mixture model and then measure the $\chi^{2}_{\rm PQM}$ value of our test, with $n_{R}=100$. The $\chi^{2}_{\rm PQM}$ value is approximately distributed as a chi-squared distribution with $n_{R}-1$ degrees of freedom, with the approximation getting better as the number of samples gets larger, as expected from § 2.1.
This test, while simple, is important, as it shows that the proposed method is not biased towards rejecting the null hypothesis. We show further null tests for complex distributions in § A, and explore approximations that can be used when a limited number of samples are available in § I.
@@ 3.2 Validation of sampling methods
[Uncaptioned image] Figure 4: Computational cost of various sample-based metrics vs data dimensionality. We show the time required to calculate the value of each metric for the experiment in Fig. 3, both as a function of the dimensionality of the data (left) and the number of samples (right).
@@
First, we compare PQMass to other sample-based metrics on a low-dimensional task where traditional sample-based metrics are known to perform well. We use the two-dimensional Gaussian mixture model used in Midgley et al. (2023), which has been used for benchmarking various sampling methods. We generate samples using three representative methods: flow-annealed importance sampling bootstrap (FAB, Midgley et al., 2023) a Markov chain Monte Carlo (Metropolis et al., 1953) algorithm as implemented in emcee (Foreman-Mackey et al., 2013) [^2] Note that MCMC does not produce independent samples, as samples within the samples are correlated. While we randomly subsample $0.01$ of the total number of samples in each chain, there could still be spurious correlations between these, affecting, the MCMC results. However, we find the MCMC results to be significantly worse than competing methods in these experiments, far beyond the possible effects of spurius correlations., and gradient-guided nested sampling (GGNS, Skilling, 2006; Lemos et al., 2023b). For each method, we generate $10,000$ samples and compare them with $10,000$ samples from the true distribution. We calculate the 2-Wasserstein distance, the mean maximum discrepancy with a radial basis function (RBF) kernel, and the Jensen-Shannon divergence (after performing a step of kernel density estimation); and compare them to the $\chi_{\rm PQM}^{2}$ value with $n_{R}=100$. For PQMass, we repeat this process $20$ times, resampling the reference points, and report the standard deviation. We show the results in Table 1. PQMass is in good agreement, and the $\chi^{2}_{\rm PQM}$ values correlate well with the other sample-based metrics, with further comparisons in § H.
@@
We then repeat this experiment on a higher-dimensional task: Neal’s 10-dimensional funnel distribution (Neal, 2003), another common sampling benchmark. PQMass is again in good agreement with MMD, but not with the Wasserstein distance. We show samples from each model in § C, which show that, visually, PQMass correctly identifies the best-performing sampling methods.
@@
Finally, we study the scaling of PQMass with the dimensionality of the data. We use a uniform mixture of $10$ Gaussians in $\mathbb{R}^{d}$, for varying $d$. For each $d$, we generate $5,000$ samples and compare them with $5,000$ samples from the same distribution. We calculate the $\chi^{2}_{\rm PQM}$ value of our test, with $n_{R}=100$, as well as other sample-based metrics. We then repeat the experiment, but comparing the true distribution with the distribution with one mode dropped. We show the results in Fig. 3. While $\mathcal{W}_{2}$ and linear MMD get increasingly noisier as the dimensionality increases, PQMass remains stable. Furthermore, the computational cost of radial basis MMD scales far worse with dimensionality than PQMass (Fig. 4). MMD with a linear kernel is the cheapest computationally, but, as shown in Fig. 3, it is too noisy to be usable even in simple high-dimensional problems. We show results with larger number of modes dropped in § F.1.
@@ 3.3 Assessing image generative models
In this section, we track the value of $\chi^{2}_{\rm PQM}$ between model-generated and ground truth samples as we train two generative models, a variational autoencoder (VAE; Kingma & Welling, 2014) and a denoising diffusion model (Ho et al., 2020; Song et al., 2021) (see § D for model details). We train on the MNIST [^3] http://yann.lecun.com/exdb/mnist/ train set and compare generated samples with the MNIST test set. After each epoch of training, we generate $2000$ samples and compare them with $2000$ test samples by computing $\chi^{2}_{\rm PQM}$ with $n_{R}=100$. We execute this experiment over the first $100$ training epochs to highlight the early stages of training for the models and repeat the experiment twice. As shown in Fig. 5, as training progresses, the metric stabilizes, with fluctuations due to the stochasticity of training and the unique samples generated at each epoch.
As shown in § E.1, the VAEs fail to capture correct structures of the images, while the diffusion-generated images are more realistic. This is reflected in the low $\chi^{2}_{\rm PQM}$ values for the diffusion model, which plateau at around $\chi^{2}_{\rm PQM}$ closer to $\approx 99$, the ideal value.
[Uncaptioned image] Figure 5: Tracking training progress of generative models on MNIST. After each training epoch, we generate samples and compute the $\chi^{2}_{\rm PQM}$ value to assess their quality against test samples. We perform this experiment twice for each model, and see that the values agree even though unique samples were generated for each run. The diffusion model can achieve values closer to ideal ($\chi^{2}_{\rm PQM}$ equal to the number of regions $n_{R}-1$, here $99$). $\chi_{\rm PQM}^{2}$ was evaluated by retessellating and resampling the data sets.
@@
To verify that our method can measure the diversity of samples from generative models, we retrain the variational autoencoder and diffusion model described above on subsets of MNIST where all images of $N$ randomly chosen classes are missing, for varying $N$. The results, averaged over random runs and samples of classes, are shown in Fig. 7. As expected, the value of $\chi^{2}_{\rm PQM}$ increases as we drop more classes for both models. In addition, we verify the ability of PQMass to detect the presence of a small mode in one of the sets in § F.2.
@@
Next, we consider the assessment of pretrained models on larger image datasets. We perform an experiment inspired by Stein et al. (2023); Jiralerspong et al. (2023), focusing on the CIFAR-10444https://www.cs.toronto.edu/~kriz/cifar.html and FFHQ256 (Karras et al., 2019) datasets. Following Stein et al. (2023), we first use the human error rate as a measure of the fidelity of the generative models. We compare generated samples from each of the models to the test data for both datasets. We repeat the comparison $20$ times, each time varying the reference points. We show the results in Fig. 6. We find that our chi-squared values visibly correlate with the human error rate, indicating that PQMass can effectively measure the fidelity and diversity of generative models.
@@
However, PQMass can also work in feature space: Table 3 shows a comparison of PQMass applied in pixel space and using InceptionV3 features (Szegedy et al., 2016).
@@
Next, we consider the detection of additive noise on ImageNet (Deng et al., 2009). We first run PQMass on a subset of the training data with added Gaussian noise of different variance, comparing it to the test data (see § E.2 for visualization). We repeat the experiment by replacing PQMass with FLD using the DinoV2 (Oquab et al., 2024) feature extractor. The results of both experiments are shown in Table 2. PQMass is more sensitive to noise, especially in the high noise regime, demonstrating that PQMass not only scales well with dimension (see § F.5 for further demonstrations of the scalability of PQMass to high-dimensional settings), but also provides a measure of distribution shift caused by increasing noise added to the data.
[…]
@@
While PQMass can measure the fidelity and diversity of model-generated samples, it does not directly measure their novelty. One approach for detecting memorization or overfitting would be to follow Jiralerspong et al. (2023) in looking at the generalization gap for PQMass, i.e., how this metric differs between comparing the generated samples to the training data and comparing the generated samples to the validation data. However, we propose a way to detect overfitting from $\chi^{2}_{\rm PQM}$ on the training data directly. Indeed, we know that a value that is too low is indicative of overfitting. For a large number of regions $n_{R}$ (i.e., $n_{R}-1$ degrees of freedom), we know the chi-square distribution is approximately Gaussian (Fig. 2). Therefore, we can use the asymmetry of this distribution to get a p-value that penalizes suspiciously low values of $\chi^{2}_{\rm PQM}$, by considering the mirror reflection of the chi-square value around the maximum:
[…]
@@
We show the results for in Fig. 8. The top left panel shows that the gap between $\chi^{2}_{\rm PQM}$ when comparing to the training data and when comparing to the validation data increases with the number of samples we replace. The top right panel shows the two-sided p-value, which decreases as the number of copied samples increases. These results show that both generalization gap and two-sided p-value can be used to detect memorization in generative models. The bottom two panels show the effectiveness of the p-value in (7) in capturing memorization and its agreement with an established metric, the authenticity score authpct (Alaa et al., 2022).
Experiments in the Appendix demonstrate the effectiveness of PQMass for other data modalities, including higher-dimensional images from scientific applications (§ F.5), time series (§ F.3), tabular data (§ F.4) and protein sequences (§ F.6).

### v3

@@ 2.1 Testing equality of quantized distributions using samples
With this definition, (1) can be restated: $p=q$ if and only if $\pi_{*}p=\pi_{*}q$ for all partitions $\mathbb{R}^{d}=R_{1}\sqcup R_{2}$, or, more generally, $\mathbb{R}^{d}=R_{1}\sqcup\dots\sqcup R_{n_{R}}$ (note that $\pi$ implicitly depends on the partition). As we will return to in §˜2.2, it is sometimes sufficient to consider restricted classes of partitions.
The main idea behind PQMass is that the probability that $p=q$ is upper-bounded by the probability that $\pi_{*}p=\pi_{*}q$ for a given partition. Next, we will consider how to test the hypothesis that $\pi_{*}p=\pi_{*}q$ using samples from $p$ and $q$ given a fixed partition. In §˜2.2 we discuss the choice of partition and propose an effective method.
@@ 2.2 Choosing an effective partition
The sets $R_{1},\dots,R_{n_{R}}$ then partition $\mathbb{R}^{d}$ into nonoverlapping regions, and the function $\pi$ that maps points to the index of the containing region is defined by $\pi(x)=\operatorname*{arg\,min}_{j}\|x-z_{j}\|$, breaking ties in favour of the lower index (see Fig.˜1, right).
While the definition (6) uses the $L^{2}$ distance metric, one can define Voronoi cells using any metric $D:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$. In Table˜3 we compare this choice to that of using the $L^{1}$ metric, as well as with the choice of using a metric defined in a feature space.
@@
For the validity of the test in §˜2.1, the choice of the regions should be independent of the points used to perform the test. Thus one can use a small subset of the available sample points as reference points to construct a tessellation, then evaluate the $\chi^{2}_{\rm PQM}$ statistic using the remaining points.
@@ 2.3 Consistency guarantees: What information is lost by PQMass?
Since $\pi$ depends on $r$, the probability referred to in Prop.˜2.1 is over the choice of $n_{R}$ samples from $r$. Of course, for "bad" choices of reference points, we may be unlucky and not distinguish $p$ and $q$ (if the two resulting cells have the same mass under $p$ and $q$). The proposition states that with positive probability, this does not happen. Similarly, as the number of references points $n_{R}$ grows to infinity, the probability that the test distinguishes two distributions approaches one. This can be understood as a consequence of the fact that the Voronoi cells ‘tessellate’ the space more finely as $n_{R}$ grows.
@@
Note that the $n_{R}\to\infty$ limit in Prop.˜2.2 does not refer to performing the test with $n_{R}=\infty$" to compare two $\chi^{2}$ distributions with infinite degrees of freedom. The proposition states that the probability that the tessellation defined using $n_{R}$ random reference points sampled from $r$ separates $p$ and $q$ – which is a well-defined quantity for any $n_{R}$ – approaches 1 as $n_{R}\to\infty$. These results show that, at least for ‘well-behaved’ densities, no information is lost by the PQMass statistic. By this, we mean that if two distributions are distinct, then the test has a positive probability of distinguishing them (and the probability approaches 1 in the asymptotic limit). This contrasts with tests that approximate the distributions by a certain fixed class (e.g., tests considering only summary statistics such as the second-moment statistics used by FID). These results can likely be generalized further, which we leave for future work.
Since Prop.˜2.2 is about the sensitivity of the test in the asymptotic limit, the proposition simply tells us that the sensitivity of the test approaches 1 as the number of samples grows, but does not tell us about the rate of convergence. The usefulness of the test itself is demonstrated by the empirical evidence in the experiments presented in the next section.
@@ 3.1 Null test
We start by validating PQMass on a null test, by comparing two sets of samples that are known to be equivalent. As shown in Fig.˜2, we use a Gaussian mixture model in $100$ dimensions with $20$ components as our generative model. We then repeat the following process $2^{14}$ times: We generate a number of samples (see legend) from the Gaussian mixture model and then measure the $\chi^{2}_{\rm PQM}$ value of our test, with $n_{R}=100$. The $\chi^{2}_{\rm PQM}$ value is approximately distributed as a chi-squared distribution with $n_{R}-1$ degrees of freedom, with the approximation getting better as the number of samples gets larger, as expected from §˜2.1.
This test, while simple, is important, as it shows that the proposed method is not biased towards rejecting the null hypothesis. We show further null tests for complex distributions in §˜A, and explore approximations that can be used when a limited number of samples are available in §˜I.
@@ 3.2 Validation of sampling methods
[Uncaptioned image] Figure 4: Computational cost of various sample-based metrics vs data dimensionality. We show the time required to calculate the value of each metric for the experiment in Fig.˜3, both as a function of the dimensionality of the data (left) and the number of samples (right).
@@
First, we compare PQMass to other sample-based metrics on a low-dimensional task where traditional sample-based metrics are known to perform well. We use the two-dimensional Gaussian mixture model used in Midgley et al. (2023), which has been used for benchmarking various sampling methods. We generate samples using three representative methods: flow-annealed importance sampling bootstrap (FAB, Midgley et al., 2023) a Markov chain Monte Carlo (Metropolis et al., 1953) algorithm as implemented in emcee (Foreman-Mackey et al., 2013) [^2] Note that MCMC does not produce independent samples, as samples within the samples are correlated. While we randomly subsample $0.01$ of the total number of samples in each chain, there could still be spurious correlations between these, affecting, the MCMC results. However, we find the MCMC results to be significantly worse than competing methods in these experiments, far beyond the possible effects of spurius correlations., and gradient-guided nested sampling (GGNS, Skilling, 2006; Lemos et al., 2023b). For each method, we generate $10,000$ samples and compare them with $10,000$ samples from the true distribution. We calculate the 2-Wasserstein distance, the mean maximum discrepancy with a radial basis function (RBF) kernel, and the Jensen-Shannon divergence (after performing a step of kernel density estimation); and compare them to the $\chi_{\rm PQM}^{2}$ value with $n_{R}=100$. For PQMass, we repeat this process $20$ times, resampling the reference points, and report the standard deviation. We show the results in Table˜1. PQMass is in good agreement, and the $\chi^{2}_{\rm PQM}$ values correlate well with the other sample-based metrics, with further comparisons in §˜H.
@@
We then repeat this experiment on a higher-dimensional task: Neal’s 10-dimensional funnel distribution (Neal, 2003), another common sampling benchmark. PQMass is again in good agreement with MMD, but not with the Wasserstein distance. We show samples from each model in §˜C, which show that, visually, PQMass correctly identifies the best-performing sampling methods.
@@
Finally, we study the scaling of PQMass with the dimensionality of the data. We use a uniform mixture of $10$ Gaussians in $\mathbb{R}^{d}$, for varying $d$. For each $d$, we generate $5,000$ samples and compare them with $5,000$ samples from the same distribution. We calculate the $\chi^{2}_{\rm PQM}$ value of our test, with $n_{R}=100$, as well as other sample-based metrics. We then repeat the experiment, but comparing the true distribution with the distribution with one mode dropped. We show the results in Fig.˜3. While $\mathcal{W}_{2}$ and linear MMD get increasingly noisier as the dimensionality increases, PQMass remains stable. Furthermore, the computational cost of radial basis MMD scales far worse with dimensionality than PQMass (Fig.˜4). MMD with a linear kernel is the cheapest computationally, but, as shown in Fig.˜3, it is too noisy to be usable even in simple high-dimensional problems. We show results with larger number of modes dropped in §˜F.1.
@@ 3.3 Assessing image generative models
In this section, we track the value of $\chi^{2}_{\rm PQM}$ between model-generated and ground truth samples as we train two generative models, a variational autoencoder (VAE; Kingma & Welling, 2014) and a denoising diffusion model (Ho et al., 2020; Song et al., 2021) (see §˜D for model details). We train on the MNIST [^3] http://yann.lecun.com/exdb/mnist/ train set and compare generated samples with the MNIST test set. After each epoch of training, we generate $2000$ samples and compare them with $2000$ test samples by computing $\chi^{2}_{\rm PQM}$ with $n_{R}=100$. We execute this experiment over the first $100$ training epochs to highlight the early stages of training for the models and repeat the experiment twice. As shown in Fig.˜5, as training progresses, the metric stabilizes, with fluctuations due to the stochasticity of training and the unique samples generated at each epoch.
As shown in §˜E.1, the VAEs fail to capture correct structures of the images, while the diffusion-generated images are more realistic. This is reflected in the low $\chi^{2}_{\rm PQM}$ values for the diffusion model, which plateau at around $\chi^{2}_{\rm PQM}$ closer to $\approx 99$, the ideal value.
Figure 5: Tracking training progress of generative models on MNIST. After each training epoch, we generate samples and compute the $\chi^{2}_{\rm PQM}$ value to assess their quality against test samples. We perform this experiment twice for each model, and see that the values agree even though unique samples were generated for each run. The diffusion model can achieve values closer to ideal ($\chi^{2}_{\rm PQM}$ equal to the number of regions $n_{R}-1$, here $99$). $\chi_{\rm PQM}^{2}$ was evaluated by retessellating and resampling the data sets.
@@
To verify that our method can measure the diversity of samples from generative models, we retrain the variational autoencoder and diffusion model described above on subsets of MNIST where all images of $N$ randomly chosen classes are missing, for varying $N$. The results, averaged over random runs and samples of classes, are shown in Fig.˜7. As expected, the value of $\chi^{2}_{\rm PQM}$ increases as we drop more classes for both models. In addition, we verify the ability of PQMass to detect the presence of a small mode in one of the sets in §˜F.2.
@@
Next, we consider the assessment of pretrained models on larger image datasets. We perform an experiment inspired by Stein et al. (2023); Jiralerspong et al. (2023), focusing on the CIFAR-10444https://www.cs.toronto.edu/~kriz/cifar.html and FFHQ256 (Karras et al., 2019) datasets. Following Stein et al. (2023), we first use the human error rate as a measure of the fidelity of the generative models. We compare generated samples from each of the models to the test data for both datasets. We repeat the comparison $20$ times, each time varying the reference points. We show the results in Fig.˜6. We find that our chi-squared values visibly correlate with the human error rate, indicating that PQMass can effectively measure the fidelity and diversity of generative models.
@@
However, PQMass can also work in feature space: Table˜3 shows a comparison of PQMass applied in pixel space and using InceptionV3 features (Szegedy et al., 2016).
@@
Next, we consider the detection of additive noise on ImageNet (Deng et al., 2009). We first run PQMass on a subset of the training data with added Gaussian noise of different variance, comparing it to the test data (see §˜E.2 for visualization). We repeat the experiment by replacing PQMass with FLD using the DinoV2 (Oquab et al., 2024) feature extractor. The results of both experiments are shown in Table˜2. PQMass is more sensitive to noise, especially in the high noise regime, demonstrating that PQMass not only scales well with dimension (see §˜F.5 for further demonstrations of the scalability of PQMass to high-dimensional settings), but also provides a measure of distribution shift caused by increasing noise added to the data.
Figure 8: Novelty. PQMass (Y-axis) when a fraction of the generated samples are replaced by samples taken from the training set (X-axis). We repeat the comparison $20$ times, each time changing the reference points, and report the standard deviation. We show the value of $\chi^{2}_{\rm PQM}$ in the top left panel, for the test (red) and train (blue), and the p-value for the training data on the top right. We see that, as memorization increases, the gap between the train and test $\chi^{2}_{\rm PQM}$ increases, and the p-value goes down. The bottom pannels compare $\text{p-value}_{{\rm overfit}}(\chi^{2}_{\rm{PQM}})$ with the $C_{T}$ score and the percentage of authentic samples authpct, two metrics thats measure novelty.
@@
While PQMass can measure the fidelity and diversity of model-generated samples, it does not directly measure their novelty. One approach for detecting memorization or overfitting would be to follow Jiralerspong et al. (2023) in looking at the generalization gap for PQMass, i.e., how this metric differs between comparing the generated samples to the training data and comparing the generated samples to the validation data. However, we propose a way to detect overfitting from $\chi^{2}_{\rm PQM}$ on the training data directly. Indeed, we know that a value that is too low is indicative of overfitting. For a large number of regions $n_{R}$ (i.e., $n_{R}-1$ degrees of freedom), we know the chi-square distribution is approximately Gaussian (Fig.˜2). Therefore, we can use the asymmetry of this distribution to get a p-value that penalizes suspiciously low values of $\chi^{2}_{\rm PQM}$, by considering the mirror reflection of the chi-square value around the maximum:
$\text{p-value}_{{\rm overfit}}(\chi^{2}_{\rm{PQM}})\equiv\int_{-\infty}^{2n_{R}-\chi^{2}_{\rm PQM}}\chi^{2}_{n_{R}-1}(z)dz.$ (7)
@@
We show the results for in Fig.˜8. The top left panel shows that the gap between $\chi^{2}_{\rm PQM}$ when comparing to the training data and when comparing to the validation data increases with the number of samples we replace. The top right panel shows the two-sided p-value, which decreases as the number of copied samples increases. These results show that both generalization gap and two-sided p-value can be used to detect memorization in generative models. The bottom two panels show the effectiveness of the p-value in (7) in capturing memorization and its agreement with an established metric, the authenticity score authpct (Alaa et al., 2022).
Experiments in the Appendix demonstrate the effectiveness of PQMass for other data modalities, including higher-dimensional images from scientific applications (§˜F.5), time series (§˜F.3), tabular data (§˜F.4) and protein sequences (§˜F.6).